Floyd is an aspiring music artist. He has a record contract that pays him a base rate of 

$200

 a month and an additional 

$12

 for each album that he sells. Last month he earned a total of 

$644

Write an equation to determine the number of albums (a)

 Floyd sold last month.





1
Expert's answer
2021-12-13T17:27:56-0500

Let Floyd sells x albums in a month.

Additional income from albums at the rate $12 for each album is $12x

Base rate per month is $200.

So his income in the month is

$( 12x + 200 )

In last his income was $644

So the equation to find the number of albums is

12x + 200 = 644

Solving it we get

12x = 644 - 200

=> 12x = 444

=> x = "\frac{444}{12}=37"

So in last month Floyd sold 37 albums.